A tent is in the shape of a triangular prism. It's base is 8 feet, height is 5 feet, and is 12 feet long. How much canvas is needed to make the tent, including the bottom?
Answer by Alan3354(69443)

A tent is in the shape of a triangular prism. It's base is 8 feet, height is 5 feet, and is 12 feet long. How much canvas is needed to make the tent, including the bottom?
----------------
The bottom area = 8*12 = 96 sq ft
Assuming the 2 sides are equal:
Sides = sqrt(5^2 + 4^2) = sqrt(41)
Each side's area = 12sqrt(41) sq ft
Each end is a triangle, Area = b*h/2
Total = 96 + 24*sqrt(41) + b*h
=========================
PS it's = it is